OpenStudy (anonymous):

Explain why (–4x)0 = 1, but –4x0 = –4

OpenStudy (anonymous):

for the same reason that \[3\times 2^2=3\times 4=12\] whereas \[(3\times 2)^2=6^2=36\]

OpenStudy (anonymous):

"order of operations" in other words

OpenStudy (anonymous):

can you explain more of the problem

OpenStudy (anonymous):

\((-4x)^0=1\) because anything raised to the power of zero is, by definition, one \(-4x^0=-4\) because \(x^0=1\) and \(-4\times 1=-4\)

OpenStudy (anonymous):

Thank you this is exactly what i needed

OpenStudy (anonymous):

yw